# Vision-Language-Action Models: Concepts, Progress, Applications and Challenges

### v1

@@ 2.2 Multimodal Integration: From Isolated Pipelines to Unified Agents
A central advancement in the emergence of VLA models lies in their ability to perform multimodal integration, the joint processing of vision, language, and action within a unified architecture. Traditional robotic systems treated perception, natural language understanding, and control as discrete modules, often linked through manually defined interfaces or data transformations [109, 20, 168]. For instance, classic pipeline-based frameworks required a perception model to output symbolic labels, which were then mapped by a planner to specific actions—frequently with domain-specific hand engineering [138, 90]. These approaches lacked adaptability, failed in ambiguous or novel environments, and could not generalize instructions beyond pre-encoded templates.
@@
In contrast, modern VLAs fuse modalities end-to-end using large-scale pretrained encoders and transformer-based architectures [188]. This shift enables the model to interpret visual observations and linguistic instructions within the same computational space, allowing flexible, context-aware reasoning [99]. For example, in the task “Pick up the red ripe apple,” (Figure 5) the vision encoder—typically a Vision Transformer (ViT) or ConvNeXt—segments and classifies objects in the scene (e.g., apples, leaves, background), identifying color and ripeness attributes [187]. Meanwhile, the language model, often a variant of T5, GPT, or BERT, encodes the instruction into a high-dimensional embedding. These representations are then fused via cross-attention or joint tokenization schemes, producing a unified latent space that informs the action policy [68].
@@
This multimodal synergy was first effectively demonstrated in CLIPort [157], which used CLIP embeddings for semantic grounding and a convolutional decoder for pixel-level manipulation. CLIPort bypassed the need for explicit language parsing and directly conditioned visuomotor policies on natural language. Similarly, VIMA [86] advanced this approach by employing a transformer encoder to jointly process object-centric visual tokens and instruction tokens, enabling few-shot generalization across spatial reasoning tasks.
@@
Recent developments push this fusion further by incorporating temporal and spatial grounding. VoxPoser [78] employs voxel-level reasoning to resolve ambiguities in 3D object selection, while RT-2 [224] fuses visual-language tokens into a unified transformer that supports zero-shot generalization to unseen instructions. Another noteworthy contribution is Octo [167], which introduces a memory-augmented transformer that enables long-horizon decision-making across diverse scenes, demonstrating the scalability of joint perception-language-action learning.
@@
Crucially, VLAs offer robust solutions to challenges in real-world grounding. For example, Occllama [183] handles occluded object references through attention-based mechanisms, while ShowUI [108] demonstrates natural language interfaces that allow non-expert users to command agents through voice or typed input. These capabilities are only possible because the integration is not limited to surface-level fusion; rather, it captures semantic, spatial, and temporal alignment across modalities.
@@ 2.3 Tokenization and Representation: How VLAs Encode the World
A core innovation that sets VLA models apart from conventional vision-language architectures lies in their token-based representation framework, which enables holistic reasoning over perceptual [125, 215], linguistic, and physical action spaces [106]. Inspired by autoregressive generative models like transformers, modern VLAs encode the world using discrete tokens that unify all modalities—vision, language, state, and action into a shared embedding space [110]. This allows the model to not only understand “what needs to be done” (semantic reasoning), but also “how to do it” (control policy execution) in a fully learnable and compositional way [192, 117, 170].
@@
    For instance, as depicted in Figure 7in a task such as “stack the green blocks on the red tray,” the image of a cluttered tabletop is processed through a vision encoder like ViT or ConvNeXt, while the instruction is embedded by a large language model (e.g., T5 or LLaMA). These are then transformed into a sequence of prefix tokens that establish the model’s initial understanding of the goal and environmental layout. This shared representation enables cross-modal grounding, allowing the system to resolve spatial references (e.g., “on the left,” “next to the blue cup”) and object semantics (“green blocks”) across both modalities.
    Figure 7: The diagram illustrates the end-to-end tokenization and representation process in VLA models. Visual input (e.g., cluttered tabletop) is encoded by a vision encoder (e.g., ViT), while natural language instructions (e.g., “stack the green blocks”) are processed by a language encoder (e.g., T5). The system fuses prefix, state, and action tokens through a transformer and autoregressively predicts motor actions.
@@
    State Tokens: Embedding the Robot’s Configuration: In addition to perceiving external stimuli, VLAs must be aware of their internal physical state [186, 111]. This is achieved through the use of state tokens, which encode real-time information about the agent’s configuration—joint positions, force-torque readings, gripper status, end-effector pose, and even the locations of nearby objects [97]. These tokens are crucial for ensuring situational awareness and safety, especially during manipulation or locomotion [163, 81].
    Figure 8 illustrates how VLA models utilize state tokens to enable dynamic, context-aware decision-making in both manipulation and navigation settings. In Figure 8a, a robot arm is shown partially extended near a fragile object. In such scenarios, state tokens play a critical role by encoding real-time proprioceptive information, such as joint angles, gripper pose, and end-effector proximity. These tokens are continuously fused with visual and language-based prefix tokens, allowing the transformer to reason about physical constraints. The model can thus infer that a collision is imminent and adjust the motor commands accordingly—e.g., rerouting the arm trajectory or modulating force output. In mobile robotic platforms, as depicted in Figure 8b, state tokens encapsulate spatial features such as odometry, LiDAR scans, and inertial sensor data. These are essential for terrain-aware locomotion and obstacle avoidance. The transformer model integrates this state representation with environmental and instructional context to generate navigation actions that dynamically adapt to changing surroundings. Whether grasping objects in cluttered environments or autonomously navigating uneven terrain, state tokens provide a structured mechanism for situational awareness, enabling the autoregressive decoder to produce precise, context-informed action sequences that reflect both internal robot configuration and external sensory data.
@@
    For instance, in the apple-picking task as depicted in Figure 9, the model may receive prefix tokens that include the image of the orchard and the text instruction. The state tokens describe the robot’s current arm posture and whether the gripper is open or closed. Action tokens are then predicted step by step to guide the robotic arm toward the apple, adjust the gripper orientation, and execute a grasp with appropriate force. The beauty of this approach is that it allows transformers, which are traditionally used for text generation, to now generate sequences of physical actions in a manner similar to generating a sentence—only here, the sentence is the motion.
@@
To operationalize the VLA paradigm in robotics, we present in Figure 9 a structured pipeline that demonstrates how multimodal information—specifically vision, language, and proprioceptive state—is encoded, fused, and converted into executable action sequences. This end-to-end loop allows a robot to interpret complex tasks like “pick the ripe apple near the green leaf” and execute precise, context-sensitive manipulations. The system begins with multimodal input acquisition, where three distinct data streams are collected: visual observations (e.g., RGB-D frames), natural language commands, and real-time robot state information (e.g., joint angles or velocity). These are independently tokenized into discrete embeddings using pretrained modules [41, 212]. As depicted in the diagram, the image is processed through a Vision Transformer (ViT) backbone to generate vision tokens, the instruction is parsed by a language model such as BERT or T5 to produce language tokens, and state inputs are transformed via a lightweight MLP encoder into compact state tokens.
@@
These tokens are then fused using a cross-modal attention mechanism, where the model jointly reasons over object semantics, spatial layout, and physical constraints [61]. This fused representation forms the contextual basis for decision-making [74, 116]. In Figure 9, this is denoted as the multimodal fusion step. The fused embedding is passed into an autoregressive decoder—typically a transformer—that generates a series of action tokens. These tokens may correspond to joint displacements, gripper force modulation, or high-level motor primitives (e.g., “move to grasp pose”, “rotate wrist”). The action tokens are subsequently translated into control commands and passed to the execution loop, which closes the perception-action cycle by feeding back the robot’s updated state, thus informing the next inference step. This closed-loop mechanism enables the model to dynamically adapt to perturbations, object shifts, or occlusions in real time [206, 120, 194].
@@
To offer concrete implementation details, Algorithm 1 formalizes the VLA tokenization process. Given an RGB-D frame $I$, natural language instruction $T$, and joint angle vector $\theta$, the algorithm produces a set of action tokens that can be executed in sequence. The image $I$ is processed via a ViT to produce $V$, a set of 400 visual tokens. In parallel, the instruction $T$ is encoded by a BERT model to yield $L$, a sequence of 12 semantic language tokens. Simultaneously, robot state $\theta$ is passed through a multilayer perceptron to generate a 64-dimensional state embedding $S$. These tokens are then fused via a cross-attention module to produce a shared 512-dimensional representation $F$, capturing the semantics, intent, and situational awareness needed for grounded action. Finally, a policy decoder such as FAST [133] maps the fused features into 50 discrete action tokens, which can then be decoded into motor commands $\tau_{1:N}$.
@@
The decoding process is implemented using a transformer-based architecture, as shown in the code snippet titled Action Prediction Code. A ‘Transformer‘ object is initialized with 12 layers, a model dimension of 512, and 8 attention heads. The fused tokens are passed to the decoder, which autoregressively predicts the next most likely action token conditioned on previous tokens and context. The final motor command sequence is obtained by detokenizing the output. This implementation mirrors how text generation works in large language models, but here the “sentence” is a motion trajectory—a novel repurposing of natural language generation techniques for physical action synthesis.
@@ 2.4 Learning Paradigms: Data Sources and Training Strategies
Training VLA models requires a hybrid learning paradigm that integrates both semantic knowledge from the web and task-grounded information from robotics datasets [30]. As shown in prior sections, the multimodal architecture of VLAs must be exposed to diverse forms of data that support language understanding, visual recognition, and motor control. This is typically achieved through two primary data sources.
@@
However, semantic understanding alone is insufficient for physical task execution [36, 178, 107]. Thus, the second phase focuses on grounding the model in embodied experience [178]. Robot trajectory datasets—collected either from real-world robots or high-fidelity simulators—are used to teach the model how language and perception translate into action [54]. These include datasets like RoboNet [37], BridgeData [50], and RT-X [175], which provide video-action pairs, joint trajectories, and environment interactions under natural language instructions [123]. Demonstration data may come from kinesthetic teaching, teleoperation, or scripted policies [89, 12]. This phase typically employs supervised learning (e.g., behavior cloning) [55], reinforcement learning (RL), or imitation learning to train the autoregressive policy decoder to predict action tokens based on fused visual-language-state embeddings [65].
Recent works increasingly adopt multistage or multitask training strategies. For example, models are often pretrained on vision-language datasets using masked language modeling, then fine-tuned on robot demonstration data using token-level autoregressive loss [94, 221, 195]. Others use curriculum learning, where simpler tasks (e.g., object pushing) precede more complex ones (e.g., multistep manipulation) [217]. Some approaches further leverage domain adaptation such as in OpenVLA [94] or sim-to-real transfer to bridge the gap between synthetic and real-world distributions [96]. By unifying semantic priors with task execution data, these learning paradigms allow VLA models to generalize across tasks, domains, and embodiments—forming the backbone of scalable, instruction-following agents capable of robust real-world operation.
@@
Recent architectures, such as Google DeepMind’s RT-2 (Robotic Transformer 2) [224], have demonstrated this principle in action. RT-2 treats action generation as a form of text generation, where each action token corresponds to a discrete command in a robot’s control space. Because the model is trained on both web-scale multimodal data and thousands of robot demonstrations, it can flexibly interpret novel instructions and perform zero-shot generalization to new objects and tasks—something that was largely impossible with traditional control systems or even early multimodal models.
@@ 2.5 Adaptive Control and Real-Time Execution
Another strength of VLAs lies in their ability to perform adaptive control, using real-time feedback from sensors to adjust behavior on the fly [153]. This is particularly important in dynamic, unstructured environments like orchards, homes, or hospitals, where unexpected changes (e.g., wind moving an apple, lighting changes, human presence) can alter the task parameters. During execution, state tokens are updated in real time, reflecting sensor inputs and joint feedback [195]. The model can then revise its planned actions accordingly. For instance, in the apple-picking scenario, if the target apple shifts slightly or another apple enters the field of view, the model dynamically reinterprets the scene and adjusts the grasp trajectory. This capability mimics human-like adaptability and is a core advantage of VLA systems over pipeline-based robotics.

### v2

@@ 2.2 Multimodal Integration: From Isolated Pipelines to Unified Agents
A central advancement in the emergence of VLA models lies in their ability to perform multi-modal integration, the joint processing of vision, language, and action within a unified architecture. Traditional robotic systems treated perception, natural language understanding, and control as discrete modules, often linked through manually defined interfaces or data transformations [140, 21, 219]. For instance, classic pipeline-based frameworks required a perception model to output symbolic labels, which were then mapped by a planner to specific actions frequently with domain-specific hand engineering [178, 117]. These approaches lacked adaptability, failed in ambiguous or unseen environments, and could not generalize instructions beyond pre-defined templates.
@@
In contrast, modern VLAs fuse modalities end-to-end using large-scale pretrained encoders and transformer-based architectures [244]. This shift enables the model to interpret visual observations and linguistic instructions within the same computational space, allowing flexible, context-aware reasoning [128]. For example, in the task “Pick the ripe apples” (Figure 5), the vision encoder—typically a Vision Transformer (ViT) or ConvNeXt—parses the scene to localize and categorize relevant objects (e.g., fruit, foliage, background) and infer ripeness-related visual cues based on learned texture, shape, and contextual features rather than fixed color assumptions [243]. Meanwhile, the language model, often a variant of T5, GPT, or BERT, encodes the instruction into a high-dimensional embedding. These representations are then fused via cross-attention or joint tokenization schemes, producing a unified latent space that informs the action policy [86].
@@
This multimodal synergy was first effectively demonstrated in CLIPort [202], which takes an RGB image of a tabletop scene and a natural language instruction (e.g., “place the blue block on the red square”) as inputs, encodes them using CLIP for semantic grounding, and outputs pixel-level pick-and-place action distributions via a convolutional transport decoder. By directly conditioning visuomotor policies on language embeddings, CLIPort eliminates explicit language parsing and enables end-to-end language-conditioned manipulation. Similarly, VIMA [112] advanced this approach by employing a transformer encoder to jointly process object-centric visual tokens and instruction tokens, enabling few-shot generalization across spatial reasoning tasks.
@@
Recent developments push this fusion further by incorporating temporal and spatial grounding. VoxPoser [100] employs voxel-level reasoning to resolve ambiguities in 3D object selection by composing pretrained vision-language models and classical motion planners, notably achieving zero-shot manipulation without task-specific training data. In contrast, RT-2 [299] fuses visual-language tokens and action representations within a unified transformer, co-trained on large-scale internet vision-language corpora and over 100,000 real-robot demonstrations from the RT-1 dataset, enabling zero-shot generalization to unseen instructions. Another noteworthy contribution is Octo [218], which introduces a memory-augmented transformer trained on more than four million robot trajectories collected across diverse robots and environments via the Open X-Embodiment dataset, supporting long-horizon decision-making and demonstrating the scalability of joint perception–language–action learning.
@@
Crucially, VLAs offer robust solutions to challenges faced in real-world grounding. For example, Occllama [239] handles occluded object references through attention-based mechanisms, while ShowUI [139] demonstrates natural language interfaces that allow non-expert users to command agents through voice or typed input. These capabilities are only possible because the integration is not limited to surface-level fusion; rather, it captures semantic, spatial, and temporal alignment across modalities.
@@ 2.3 Tokenization and Representation: How VLAs Encode the World
A core innovation that sets VLA models apart from conventional vision-language architectures lies in their token-based representation framework, which enables holistic reasoning over perceptual [161, 286], linguistic, and physical action spaces [136]. Inspired by autoregressive generative models like transformers, modern VLAs encode the world using discrete tokens that unify all modalities vision, language, state, and action into a shared embedding space [142]. This allows the model to not only understand “what needs to be done” (semantic reasoning), but also “how to do it” (control policy execution) in a fully learnable and compositional way [248, 150, 221].
@@
    For instance, as depicted in Figure 7 in a task such as “stack the green blocks on the red tray,” the image of a cluttered tabletop is processed through a vision encoder like ViT or ConvNeXt, while the instruction is embedded by an LLM (e.g., T5 or LLaMA). These are then transformed into a sequence of prefix tokens that establish the model’s initial understanding of the goal and environmental layout. This shared representation enables cross-modal grounding, allowing the system to resolve spatial references (e.g., “on the left,” “next to the blue cup”) and object semantics (“green blocks”) across both modalities.
    Figure 7: A diagram illustrating the end-to-end tokenization and representation process in VLA models. Visual input (e.g., cluttered tabletop) is encoded by a vision encoder (e.g., ViT), while natural language instructions (e.g., “stack the green blocks”) are processed by a language encoder (e.g., T5). The system fuses prefix, state, and action tokens through a transformer and autoregressively predicts motor actions.
@@
    State Tokens: Embedding the Robot’s Configuration: In addition to perceiving external stimuli, VLAs must be aware of their internal physical state [242, 143]. This is achieved through the use of state tokens, which encode real-time information about the agent’s configuration joint positions, force-torque readings, gripper status, end-effector pose, and even the locations of nearby objects [126]. These tokens are crucial for ensuring situational awareness and safety, especially during manipulation or locomotion [211, 105].
    Figure 8 illustrates how VLA models utilize state tokens to enable dynamic, context-aware decision-making in both manipulation and navigation settings. In Figure 8a, a robot arm is shown partially extended near a fragile object. In such scenarios, state tokens play a critical role by encoding real-time proprioceptive information, such as joint angles, gripper pose, and end-effector proximity. These tokens are continuously fused with visual and language-based prefix tokens, allowing the transformer to reason about physical constraints. The model can thus infer that a collision is imminent and adjust the motor commands accordingly e.g., rerouting the arm trajectory or modulating force output. In mobile robotic platforms, as depicted in Figure 8b, state tokens encapsulate spatial features such as odometry, LiDAR scans, and inertial sensor data. These are essential for terrain-aware locomotion and obstacle avoidance. The transformer model integrates this state representation with environmental and instructional context to generate navigation actions that dynamically adapt to changing surroundings. Whether grasping objects in cluttered environments or autonomously navigating uneven terrain, state tokens provide a structured mechanism for situational awareness, enabling the autoregressive decoder to produce precise, context-informed action sequences that reflect both internal robot configuration and external sensory data.
@@
    For instance, in the apple-picking task as depicted in Figure 9, the model may receive prefix tokens that include the image of the orchard and the text instruction. The state tokens describe the robot’s current arm posture and whether the gripper is open or closed. Action tokens are then predicted step by step to guide the robotic arm toward the apple, adjust the gripper orientation, and execute a grasp with appropriate force. The beauty of this approach is that it allows transformers, which are traditionally used for text generation, to now generate sequences of physical actions in a manner similar to generating a sentence only here, the sentence is the motion.
@@
To operationalize the VLA paradigm in robotics, we present in Figure 9 a structured pipeline that demonstrates how multimodal information specifically vision, language, and proprioceptive state is encoded, fused, and converted into executable action sequences. This end-to-end loop allows a robot to interpret complex tasks like “pick the ripe apple near the green leaf” and execute precise, context-sensitive manipulations. The system begins with multimodal input acquisition, where three distinct data streams are collected: visual observations (e.g., RGB-D frames), natural language commands, and real-time robot state information (e.g., joint angles or velocity). These are independently tokenized into discrete embeddings using pretrained modules [51, 282]. As depicted in the diagram, the image is processed through a Vision Transformer (ViT) backbone to generate vision tokens, the instruction is parsed by a language model such as BERT or T5 to produce language tokens, and state inputs are transformed via a lightweight MLP encoder into compact state tokens.
@@
These tokens are then fused using a cross-modal attention mechanism, where the model jointly reasons over object semantics, spatial layout, and physical constraints [76]. This fused representation forms the contextual basis for decision-making [96, 149]. In Figure 9, this is denoted as the multi-modal fusion step. The fused embedding is passed into an autoregressive decoder typically a transformer that generates a series of action tokens. These tokens may correspond to joint displacements, gripper force modulation, or high-level motor primitives (e.g., “move to grasp pose”, “rotate wrist”). The predicted action tokens are subsequently translated into low-level control commands and executed by an external, hardware-dependent execution loop, which interfaces with the robot controller to close the perception-action cycle by feeding back updated state observations for the next VLA inference step. This closed-loop mechanism enables the model to dynamically adapt to perturbations, object shifts, or occlusions in real time [278, 155, 251].
@@
To provide clear and specific implementation details, Algorithm 1 formalizes the VLA tokenization process. Given an RGB-D frame $I$, natural language instruction $T$, and joint angle vector $\theta$, the algorithm produces a set of action tokens that can be executed in sequence. The image $I$ is processed via a ViT to produce $V$, a set of 400 visual tokens. In parallel, the instruction $T$ is encoded by a BERT model to yield $L$, a sequence of 12 semantic language tokens. Simultaneously, the robot state $\theta$—including joint angles, end-effector pose, and proprioceptive signals—is encoded by a multilayer perceptron into a compact 64-dimensional state embedding $S$, providing the model with real-time awareness of the robot’s configuration and physical constraints during action generation. These tokens are then fused via a cross-attention module to produce a shared 512-dimensional representation $F$, capturing the semantics, intent, and situational awareness needed for grounded action. Finally, a policy decoder such as FAST [171] maps the fused features into 50 discrete action tokens, which can then be decoded into motor commands $\tau_{1:N}$.
@@
The decoding process is implemented using a transformer-based architecture, as shown in the code snippet titled Action Prediction Code. A Transformer decoder is instantiated with 12 layers, a model dimension of 512, and 8 attention heads. The fused multimodal tokens are provided as context, and the decoder autoregressively predicts action tokens one step at a time, where each predicted token represents the next control decision conditioned on the full multimodal context and all previously generated actions. The resulting action-token sequence is then detokenized into a continuous motor command trajectory for execution. This implementation mirrors how text generation works in LLMs, but here the “sentence” is a motion trajectory a novel repurposing of natural language generation techniques for physical action synthesis.
@@ 2.4 Learning Paradigms: Data Sources and Training Strategies
Training VLA models requires a hybrid learning paradigm that integrates both semantic knowledge from the web and task-grounded information from robotics datasets [35]. As shown in prior sections, the multi-modal architecture of VLAs must be exposed to diverse forms of data that support language understanding, visual recognition, and motor control. This is typically achieved through two primary data sources.
@@
However, semantic understanding alone is insufficient for physical task execution [44, 231, 137]. Thus, the second phase focuses on grounding the model in embodied experience [231]. Robot trajectory datasets collected either from real-world robots or high-fidelity simulators are used to teach the model how language and perception translate into action [67]. These include datasets like RoboNet [45], BridgeData [61], and RT-X [227], which provide video-action pairs, joint trajectories, and environment interactions under natural language instructions [159]. Demonstration data may come from kinesthetic teaching, teleoperation, or scripted policies [115, 13]. This phase typically employs supervised learning (e.g., behavior cloning) [68], reinforcement learning (RL), or imitation learning to train the autoregressive policy decoder to predict action tokens based on fused visual-language-state embeddings [83].
Recent works increasingly adopt multistage or multitask training strategies. For example, models are often pretrained on vision-language datasets using masked language modeling, then fine-tuned on robot demonstration data using token-level autoregressive loss [122, 295, 252]. Others use curriculum learning, where simpler tasks (e.g., object pushing) precede more complex ones (e.g., multistep manipulation) [288]. Some approaches further leverage domain adaptation such as in OpenVLA [122] or sim-to-real transfer to bridge the gap between synthetic and real-world distributions [125]. By unifying semantic priors with task execution data, these learning paradigms allow VLA models to generalize across tasks, domains, and embodiments forming the backbone of scalable, instruction-following agents capable of robust real-world operation.
@@
Recent architectures, such as Google DeepMind’s RT-2 (Robotic Transformer 2) [299], have demonstrated this principle in action. RT-2 treats action generation as a form of text generation, where each action token corresponds to a discrete command in a robot’s control space. Because the model is trained on both web-scale multi-modal data and thousands of robot demonstrations, it can flexibly interpret novel instructions and perform zero-shot generalization to new objects and tasks something that was largely impossible with traditional control systems or even with early multi-modal models.
@@ 2.5 Adaptive Control and Real-Time Execution
Another strength of VLAs lies in their ability to perform adaptive control, using real-time feedback from sensors to adjust behavior on the fly [195]. This is particularly important in dynamic, unstructured environments like orchards, homes, or hospitals, where unexpected changes (e.g., wind moving an apple, lighting changes, human presence) can alter the task parameters. During execution, state tokens are updated in real-time, reflecting sensor inputs and joint feedback [252]. The model can then revise its planned actions accordingly. For instance, in the apple-picking scenario, if the target apple shifts slightly or another apple enters the field of view, the model dynamically reinterprets the scene and adjusts the grasp trajectory. This capability mimics human-like adaptability and is a core advantage of VLA systems over pipeline-based robotics.